OpenStudy (anonymous):

From the word 'SAVIOR' Determine how many letter arrangements are possible given the following conditions. ____a. all six letter are to be used with no restriction ____b.all vowels and consonants are together. ____c.all six letter are to used S in the first position and R in the last position

OpenStudy (anonymous):

first one is \[6!=6\times 5\times 4\times 3\times2\]

OpenStudy (anonymous):

second one is \[3!\times 3!\times 2\]

OpenStudy (anonymous):

three consonants first has 6 possibilities and three vowels second has 6 possibilities, and same for three consonants second and three vowels first

OpenStudy (anonymous):

S in the first, R in the last leave 4! possibilities